OpenStudy (anonymous):

k = (Squared root 2k-4) + 2. solve for k

OpenStudy (mathstudent55):

\(k = \sqrt{2k - 4} + 2\) Is this the problem?

OpenStudy (anonymous):

yup. so would it be k-2 = squared root 2k-4??

OpenStudy (mathstudent55):

You need to have the square root alone on one side, so that when you square both sides you get rid of the square root.

OpenStudy (mathstudent55):

@akausen You were correct. Subtract 2 from both sides. @hs_beauty You are correct.

OpenStudy (anonymous):

(k-2)^2 = 2k-4

OpenStudy (anonymous):

Will I always have to use the quadratic formula??

OpenStudy (mathstudent55):

We're not ready for the quadratic formula yet, but soon. \((k - 2)^2 = (\sqrt{2k - 4} )^2\) \((k - 2)^2 = 2k - 4\) Now calculate the square of the binomial on the left side.

OpenStudy (anonymous):

would it be k^2 - 4??

OpenStudy (mathstudent55):

No. \((k - 2)^2 = 2k - 4\) \((k - 2)(k - 2) = 2k - 4\) You can rewrite the square of the binomial as I did above. Now use FOIL.

OpenStudy (anonymous):

k^2-4k+4?

OpenStudy (mathstudent55):